OpenStudy (anonymous):

differentiate log(x(root(x+1))

OpenStudy (anonymous):

\[\ln x \sqrt{x+1} =y\]

OpenStudy (watchmath):

rewrite first as \(y=\ln(x)+\frac{1}{2}\ln(x+1)\) Therefore \(y'=\frac{1}{x}+\frac{1}{2(x+1)}\)

OpenStudy (phi):

\[\ln (x \sqrt{x+1}) =y\] ?

OpenStudy (anonymous):

Can you restate your question please, I will show you a good way to deal with derivatives well at least that I use. Just split the equation into pieces and derive the pieces then place them all together

OpenStudy (anonymous):

@BlingBlong is it the way watchmath has shown thanks@ watchmath!